OpenStudy (anonymous):

just need to know what the u term is. i know how to evaluate after that. int((7^sqrt(x))*ln(7))/2*sqrt(x)) x: 4,2

OpenStudy (anonymous):

lana, can you use equation to write your question?

OpenStudy (mertsj):

\[\int\limits_{2}^{4}7\sqrt{x}(\frac{\ln 7}{2\sqrt{x}})\]

OpenStudy (mertsj):

Is that the question?

OpenStudy (anonymous):

Yeah, that's the question

OpenStudy (mertsj):

\[\int\limits_{2}^{4}7^{\sqrt{x}}(\frac{\ln 7}{2\sqrt{2})}\]

OpenStudy (mertsj):

Or is it that?

OpenStudy (mertsj):

Whoops. Make that an x in the denominator

OpenStudy (anonymous):

I don't think the asker know which one is correct!

OpenStudy (anonymous):

the 1st one is correct

OpenStudy (mertsj):

If the first one is correct, cancel out the sqrtx and you are left with only constants.

OpenStudy (anonymous):

ok, thanks

OpenStudy (anonymous):

= (14 ln7/9) [ 4^(9/14 - 2 ^(9/14) ] + C

OpenStudy (anonymous):

I thought it is x^(1/7) and x^(1/2) ?

OpenStudy (anonymous):

So confused!